# Josephson effect in a Fibonacci quasicrystal

### v1

@@ IV.2 Critical current versus junction length
The previous Sec. IV.1 demonstrated a non-trivial behavior in the critical current for different approximants and thus different junction lengths. Generally speaking, the critical current decays with junction length, related to that the Josephson effect is a mesoscopic effect relying on phase coherence being mediated across the junction, via finite wave function overlap between the two superconducting leads [154]. In this subsection, we study how this decay is influenced by quasiperiodicity from $L\sim a_{0}\ll\xi_{0}$ to $L=1000a_{0}\gg\xi_{0}$, thus varying from the short to long junction limit. We start with the case of the repeated Fibonacci approximants $C_{n}$ where $L=a_{0}N\times F_{N}$, followed by the non-repeated Fibonacci chain with more general $L$ modeled by the characteristic function Eq. (2). To avoid confusion, we point out that $L$ counts the number of bonds, while the number of sites is $L/a_{0}+1$.
@@
Beginning with a junction corresponding to a repeated Fibonacci approximant, Fig. 10 shows ${I_{\mathrm{c}}}$ as a function of $L$ for the approximant $C_{9}$ at fixed $\rho=0.8$ (a) and at fixed ${t_{\mathrm{int}}}=0.8{t_{\mathrm{B}}}$ (b). The critical current shows an overall decrease with $L$ as expected, but is superposed with a staggered behavior due to the microscopic even-odd effect described in Sec. III.2. Specifically, odd number of sites (even repetitions) leads to a robust zero-energy state which significantly increases the critical current, while for even number of sites there is only a zero-energy state at perfect Andreev reflection [141], i.e. here at ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})=\rho\sqrt{{\rho_{\mathrm{S}}}}$. Thus, the scenario ${t_{\mathrm{int}}}/{t_{\mathrm{B}}}=\rho=0.8$ at ${\rho_{\mathrm{S}}}=1$ in Fig. 10(a) signifies the maximum possible critical current, yielding an envelope (purple dash-dotted line) of all other curves. In (b) the curve for ${t_{\mathrm{int}}}/{t_{\mathrm{B}}}=\rho=0.8$ is only the maximum at even number of sites since the system with odd number of sites generally favors larger $\rho>{t_{\mathrm{int}}}/{t_{\mathrm{B}}}$ [see Fig. 9(a)]. We verify that other Fibonacci approximants follow a similar staggered behavior with an envelope curve given by the corresponding ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ expression given by Table 1. The exception to the behavior is the Fibonacci approximants which always have an odd number of sites (i.e. every third approximant), which we demonstrate in Fig. 11 for the approximant $C_{8}$. This approximant thus lacks the even-odd staggering, only illustrating the overall decay of ${I_{\mathrm{c}}}$ with increased $L$. The current is maximal for $\rho=1$ and ${t_{\mathrm{int}}}={t_{\mathrm{B}}}$ where there is a degenerate zero-energy ABS, while any other $\rho\neq 1$ splits the ABS to finite energies thus reducing the current (see Fig. 6). Interestingly, by comparing the curves in Fig. 11(b) we find for $\rho\neq 1$ a rapid decay that is approximately exponential and corresponds to the contribution from the finite-energy modes. For $\rho=1$, this is superposed with a long $\propto 1/L$ tail due to the zero-energy ABS, which is a well-known behavior in crystalline junctions [142].
@@
Next, we study the critical current decay in the more general quasiperiodic junctions without repetitions and of arbitrary length $L$, by adding one site at a time following the characteristic function in Eq. (2). We note that the crystalline junction in this case obtains an extreme oscillation due to the even-odd effect. To improve visibility, we therefore plot the crystalline results for even and odd number of sites separately. We also note that each Fibonacci approximant $C_{n}$ is represented exactly once, i.e. when $L=a_{0}F_{n}$, and we verify that this reproduces the result above. Figure 12(a) shows ${I_{\mathrm{c}}}$ as function of $L$ for a crystalline (quasiperiodic) junction as a dashed (solid) line, where Figs. 12(b) and (c) are zooms of (a) at $L<200a_{0}$ and $L>200a_{0}$, respectively. We begin by focusing on the crystalline results $\rho=1$, where even (odd) number of sites is shown as a black (gray) dashed line, again showing the well-known monotonic decay [142]. The junction with odd number of sites has a significantly higher critical current since it is an idealized scenario with an exact zero-energy state as $\Delta\theta\to\pi$. The junction with even number of sites approaches this curve as ${t_{\mathrm{int}}}\to{t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, due to also approaching an exact zero-energy state.
@@
In the quasiperiodic junction $\rho\neq 1$, Fig. 12 also illustrates an overall decay in ${I_{\mathrm{c}}}$ with $L$, but the microscopic oscillations are much less trivial, going beyond just an even-odd effect. We find that these oscillations are not random or erratic, however, but rather correlate with how the junction varies between different expressions ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ with $L$. We illustrate this Fig. 13 by plotting ${I_{\mathrm{c}}}$ as a function of $L$ with data points colored according to their functional form of ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ (as given by Table 2 in Appendix A). The figure thus shows that all data points with ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ being e.g. $\rho\sqrt{{\rho_{\mathrm{S}}}}$, $\sqrt{\rho}\sqrt{{\rho_{\mathrm{S}}}}$ or $\sqrt{{\rho_{\mathrm{S}}}}$ each follow their respective monotonic decay as indicated by the dashed lines, except for the purple curve (see discussion further below). The data points without dashed lines follow a similar trend, but these points are fewer and farther between, and we therefore omit plotting their lines for visibility reasons. Comparing the curves for the different hopping parameters in (a) and (b), the lines are shifted with respect to each other since ${I_{\mathrm{c}}}$ is maximal at ${t_{\mathrm{int}}^{*}}$, e.g. the red line has higher ${I_{\mathrm{c}}}$ than the blue one in (a) since ${t_{\mathrm{int}}}$ is closer to $\rho\sqrt{{\rho_{\mathrm{S}}}}$ than $\sqrt{{\rho_{\mathrm{S}}}}$, and vice versa in (b). Next, we comment on the spatial shape of the oscillations. Interestingly, as $L$ increases, the junction changes functional form of ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ in a quasiperiodic manner as discussed in Sec. IV.1 and Appendix A, such that the critical current ${I_{\mathrm{c}}}(L)$ therefore also varies in a quasiperiodic way. This conclusion is based on studying the first 200 values of $L$ in the characteristic function, and the first 18 approximants $C_{n}$. The exception to this behavior are the purple data points which instead show oscillations around a monotonic decay. These data points correspond to when quasiperiodicity breaks the zero-energy degeneracy, and we attribute the oscillations to variations in the size of the finite-energy shifts. We propose that studying the Andreev reflection amplitudes (i.e. with a scattering-matrix approach) might yield additional insight into these oscillations, but leave it as a future outlook.
@@
Figure 13: Same as Fig. 12(b) but for the quasiperiodic junction at $\rho=0.8$, ${t_{\mathrm{int}}}=0.8{t_{\mathrm{B}}}$ (a), and at $\rho=1.2$, ${t_{\mathrm{int}}}=0.8{t_{\mathrm{B}}}$ (b). Data points (markers) are grouped by even or odd number of sites, and by the condition for zero-energy ABS, as given by Table 2 in Appendix A. Dashes: guide to the eye showing the trend ${I_{\mathrm{c}}}(L)$ for the data points within the most populated groups.
@@
Finally, we briefly comment on how the overall magnitude of ${I_{\mathrm{c}}}$ compares between the quasiperiodic and crystalline junctions, when varying ${t_{\mathrm{int}}}\in[0.1,1.0]{t_{\mathrm{B}}}$, $\rho\in[0.1,1.4]$ and $L\in[2,1000]a_{0}$. If taking the idealized scenario for odd number of sites into account, we find that the crystalline junction quite generally shows a higher critical current except in a few cases, since the crystalline junction is ideal with perfectly ballistic extended states and exactly degenerate ABS. However, if this idealized scenario with odd number of sites is not taken into account (in either the quasiperiodic or crystalline junctions), then the quasiperiodic junction can host a significantly higher critical current by several factors, especially at low ${t_{\mathrm{int}}}$.
V Gap structure and topological invariant
@@
So far, we have considered a system at half filling through fixed chemical potential $\mu=0$ and phason angle $\phi$, at which the system behaves as a hybrid SNS Josephson junction. This also means that we have not directly probed the influence of the topological gaps or winding states on the Josephson effect. We next study this influence and demonstrate how the winding of the subgap states in the different gaps lead to distinct oscillations in the critical current ${I_{\mathrm{c}}}$, such that each winding number (i.e. gap label) can be determined, thus effectively measuring the topological invariant. This occurs when the Fermi level is located inside or close to the respective gap (i.e. by an applied gate voltage). Then by varying $\phi$, we find that ${I_{\mathrm{c}}}$ oscillates. To better describe the results, we first demonstrate in Sec. V.1 how ${I_{\mathrm{c}}}$ behaves when varying the $\mu$ at fixed $\phi$, leading to a controlled SNS to SIS transition as the Fermi level enters the gaps of the Fibonacci chain. Then, we show in Sec. V.2 that by varying the $\phi$, the winding of the topological subgap states leads to an analogous oscillation in the critical current. Finally, we extract the winding number by computing the power spectrum.
V.1 Quasiperiodic SNS to SIS transitions
To better describe the subsequent results where both $\mu$ and $\phi$ are varied, we first briefly comment on the behavior of critical current as function of $\mu$ at fixed $\phi=\pi\tau^{-1}$. The variation of $\mu$ is envisioned by the application of a gate voltage to the non-superconducting part of the junction.
@@ V.1 Quasiperiodic SNS to SIS transitions
Next, we focus on the quasiperiodic scenario $\rho\neq 1$ and note the same features with peaks and overall larger current close to half-filling (but lower bandwidth $|\mu|<1.8{t_{\mathrm{B}}}$). Importantly, however, the quasiperiodic gap structure is directly probed by a sudden drop in current as the Fermi level enters the quasiperiodic gaps ${\Delta_{\mathrm{Q}}}$. We find that the size of these gaps relative to the proximitized superconducting gap $|\Delta|$ determines the junction behavior. In particular, inside the larger gaps (${\Delta_{\mathrm{Q}}}>|\Delta|$) we find SIS behavior as that the current can become orders of magnitude smaller than outside the gaps where there is instead SNS behavior. Inside the smaller gaps (${\Delta_{\mathrm{Q}}}<|\Delta|$) the behavior more strongly depends on the junction length. For instance, in the shorter junction in Fig. 14(a) the current is still the same order of magnitude as outside the gaps due to pronounced ABS and supergap contributions, and thus of SNS type, but the gaps still cause a reduction or even complete absence of some peaks that are otherwise present in the crystalline junction. In contrast, in the longer junction in Fig. 14(b), the current is more strongly suppressed in the small gaps, showing more SIS-type behavior. Thus, the gap ratio ${\Delta_{\mathrm{Q}}}/|\Delta|$ together with the Fibonacci chain length can qualitatively alter the junction properties, and the SNS to SIS transitions in $\mu$ are fractal since the Fibonacci spectrum itself is fractal [17]. We note that the gap ratio also becomes important for emergent topological [73] and intrinsic superconductivity [70].
@@
Finally, we briefly comment on the overall magnitude of ${I_{\mathrm{c}}}$ in quasiperiodic versus crystalline junctions when varying ${t_{\mathrm{int}}}\in[0.1,1.0]{t_{\mathrm{B}}}$, $\rho\in[0.1,1.4]$ and $\mu$ between the band edges. Similar to Sec. IV.2, we find that the crystalline junction in most cases has a higher critical current due to the idealized perfect ballistic model, with exact zero-energy state in the crystalline junction. Only in a few parts of the parameter space does the quasiperiodic junction provide a significantly higher ${I_{\mathrm{c}}}$, usually when ${t_{\mathrm{int}}}$ is closer to the quasiperiodically modified expression for ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, or when ${t_{\mathrm{int}}}$ is small and $\rho$ large.
@@
V.2 Topological invariant from critical current
In the previous Sec. V.1 we investigated the behavior of $\mu$ on ${I_{\mathrm{c}}}$, demonstrating that the junction can change between SNS and SIS behavior whenever the Fermi level enter the largest topological gaps of the Fibonacci chain. We next show how the critical current can measure the winding number in each of these gaps. We begin by describing the principle idea behind connecting the winding number to the critical current, followed by our results and proof-of-principle.

### v2

@@ IV.2 Critical current versus junction length
The previous Sec. IV.1 demonstrated a non-trivial behavior in the critical current for different approximants and thus different junction lengths. Generally speaking, the critical current decays with junction length since the Josephson effect is a mesoscopic effect relying on phase coherence being mediated across the junction, via finite wave function overlap between the two superconducting leads [147]. In this subsection, we study how this decay is influenced by quasiperiodicity from $L\sim a_{0}\ll\xi_{0}$ to $L=1000a_{0}\gg\xi_{0}$, thus varying from the short to long junction limit. We start with the case of the repeated Fibonacci approximants $C_{n}$ with $L=a_{0}N\times F_{N}$ (i.e. relevant for the crystal approximants [22]), followed by the non-repeated Fibonacci chain with more general $L$ modeled by the characteristic function Eq. (2) (i.e. more relevant for the Fibonacci quasicrystals [17]). Although modern experimental techniques allows the fabrication of materials with atomic precision [98, 99, 100, 101, 102, 103, 104, 105, 106, 107], the main purpose of this section is to theoretically establish how quasiperiodicity influences the overall trend ${I_{\mathrm{c}}}(L)$.
@@
Beginning with a junction corresponding to a repeated Fibonacci approximant, Fig. 8 shows ${I_{\mathrm{c}}}$ as a function of $L$ for the approximant $C_{9}$ at fixed $\rho=0.8$ (a) and at fixed ${t_{\mathrm{int}}}=0.8{t_{\mathrm{B}}}$ (b). The critical current shows an overall decrease with $L$ as expected, but is superposed with a staggered behavior due to the microscopic even-odd effect described in Appendix A. Specifically, odd number of sites (even repetitions) leads to a robust zero-energy state which significantly increases the critical current, while for even number of sites there is only a zero-energy state at perfect Andreev reflection [146], i.e. here at ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})=\rho\sqrt{{\rho_{\mathrm{S}}}}$. Thus, the scenario ${t_{\mathrm{int}}}/{t_{\mathrm{B}}}=\rho=0.8$ at ${\rho_{\mathrm{S}}}=1$ in Fig. 8(a) signifies the maximum possible critical current, yielding an envelope (purple dash-dotted line) of all other curves. In (b) the curve for ${t_{\mathrm{int}}}/{t_{\mathrm{B}}}=\rho=0.8$ is only the maximum at even number of sites since the system with odd number of sites generally favors larger $\rho>{t_{\mathrm{int}}}/{t_{\mathrm{B}}}$ [see Fig. 7(a)]. We verify that other Fibonacci approximants follow a similar staggered behavior with an envelope curve given by the corresponding ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ expression given by Table 1. The exception to the behavior is the Fibonacci approximants which always have an odd number of sites (i.e. every third approximant), which we demonstrate in Fig. 9 for the approximant $C_{8}$. This approximant thus lacks the even-odd staggering, only illustrating the overall decay of ${I_{\mathrm{c}}}$ with increased $L$. The current is maximal for $\rho=1$ and ${t_{\mathrm{int}}}={t_{\mathrm{B}}}$ where there is a degenerate zero-energy ABS, while any other $\rho\neq 1$ splits the ABS to finite energies thus reducing the current (see Fig. 15). Interestingly, by comparing the curves in Fig. 9(b) we find for $\rho\neq 1$ a rapid decay that is approximately exponential and corresponds to the contribution from the finite-energy modes. For $\rho=1$, this is superposed with a long $\propto 1/L$ tail due to the zero-energy ABS, which is a well-known behavior in crystalline junctions [147, 146, 148, 149].
@@
Next, we study the critical current decay in the more general quasiperiodic junctions without repetitions and of arbitrary length $L$, by adding one site at a time following the characteristic function in Eq. (2). We note that the crystalline junction in this case obtains an extreme oscillation due to the even-odd effect. To improve visibility, we therefore plot the crystalline results for even and odd number of sites separately. We also note that each Fibonacci approximant $C_{n}$ is represented exactly once, i.e. when $L=a_{0}F_{n}$, and we verify that this reproduces the result above. Figure 10(a) shows ${I_{\mathrm{c}}}$ as function of $L$ for a crystalline (quasiperiodic) junction as a dashed (solid) line, where Figs. 10(b) and (c) are zooms of (a) at $L<200a_{0}$ and $L>200a_{0}$, respectively. We begin by focusing on the crystalline results $\rho=1$, where even (odd) number of sites is shown as a black (gray) dashed line, again showing the well-known monotonic decay [147, 146, 148, 149]. The junction with odd number of sites has a significantly higher critical current since it is an idealized scenario with an exact zero-energy state as $\Delta\theta\to\pi$. The junction with even number of sites approaches this curve as ${t_{\mathrm{int}}}\to{t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, due to also obtaining a zero-energy state.
@@
Figure 11: Same as Fig. 10(b) but for the quasiperiodic junction at $\rho=0.8$, ${t_{\mathrm{int}}}=0.8{t_{\mathrm{B}}}$ (a), and at $\rho=1.2$, ${t_{\mathrm{int}}}=0.8{t_{\mathrm{B}}}$ (b). Data points (markers) are grouped by even or odd number of sites, and by the condition for zero-energy ABS, as given by Table 2 in Appendix B. Dashes: guide to the eye showing the trend ${I_{\mathrm{c}}}(L)$ for the data points within the most populated groups.
In the quasiperiodic junction $\rho\neq 1$, Fig. 10 also illustrates an overall decay in ${I_{\mathrm{c}}}$ with $L$, but the microscopic oscillations are much less trivial, going beyond just an even-odd effect. We find that these oscillations are not random or erratic, however, but rather correlate with how the junction varies between different expressions ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ with $L$. We illustrate this Fig. 11 by plotting ${I_{\mathrm{c}}}$ as a function of $L$ with data points colored according to their functional form of ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ (as given by Table 2 in Appendix B). The figure thus shows that all data points with ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ being e.g. $\rho\sqrt{{\rho_{\mathrm{S}}}}$, $\sqrt{\rho}\sqrt{{\rho_{\mathrm{S}}}}$ or $\sqrt{{\rho_{\mathrm{S}}}}$ each follow their respective monotonic decay as indicated by the dashed lines, except for the purple curve (see discussion further below). We have fitted these monotonic decays and find that they are similar to the well-known exponential and power law decays in the crystalline junctions, depending on system and parameter regimes [147, 146, 148, 149]. The data points without dashed lines follow a similar trend, but these points are fewer and farther between, and we therefore omit plotting their lines for visibility reasons. Comparing the curves for the different hopping parameters in (a) and (b), the lines are shifted with respect to each other since ${I_{\mathrm{c}}}$ is maximal at ${t_{\mathrm{int}}^{*}}$, e.g. the red line has higher ${I_{\mathrm{c}}}$ than the blue one in (a) since ${t_{\mathrm{int}}}$ is closer to $\rho\sqrt{{\rho_{\mathrm{S}}}}$ than $\sqrt{{\rho_{\mathrm{S}}}}$, and vice versa in (b). Next, we comment on the spatial shape of the microscopic oscillations. Interestingly, the junction changes functional form of ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ with $L$ in a quasiperiodic manner as discussed in Sec. IV.1 and Appendix B, such that the critical current ${I_{\mathrm{c}}}(L)$ therefore also varies in a quasiperiodic way. This conclusion is based on studying the first 200 values of $L$ in the characteristic function, and the first 18 approximants $C_{n}$. The exception to this behavior are the purple data points which instead show oscillations around a monotonic decay. These data points correspond to when quasiperiodicity breaks the zero-energy degeneracy, and we attribute the oscillations to variations in the size of the finite-energy shifts. We propose that studying the Andreev reflection amplitudes (i.e. with a scattering-matrix approach) might yield additional insight into these oscillations, but leave it as a future outlook.
@@
Finally, we briefly comment on how the overall magnitude of ${I_{\mathrm{c}}}$ compares between the quasiperiodic and crystalline junctions, when varying ${t_{\mathrm{int}}}\in[0.1,1.0]{t_{\mathrm{B}}}$, $\rho\in[0.1,1.4]$ and $L\in[2,1000]a_{0}$. If taking the idealized scenario for odd number of sites into account, we find that the crystalline junction quite generally shows a higher critical current than the quasiperiodic junction except in a few cases, since the crystalline junction is ideal with perfectly ballistic extended states and exactly degenerate ABS. However, if this idealized scenario with odd number of sites is not taken into account (in either the quasiperiodic or crystalline junctions), then the quasiperiodic junction can host a significantly higher critical current than the crystalline junction by several factors, especially at low ${t_{\mathrm{int}}}$. Overall, we find microscopic quasiperiodic oscillations present in ${I_{\mathrm{c}}}(L)$, indicating significant sample-to-sample fluctuations in an experimental setup, while beyond these oscillations the current shows the same overall decay as in the crystalline scenario.
V Fractal gap structure and topological invariant
@@
So far, we have considered a system at half filling through fixed chemical potential $\mu=0$ and with fixed phason angle $\phi=\pi\tau^{-1}$, at which the system behaves as a hybrid SNS Josephson junction. The purpose of this section is to investigate how the Josephson effect is influenced by quasiperiodicity when varying $\mu$, i.e. varying an applied gate voltage in an experimental setup, as well as by the topological gaps and winding states of the Fibonacci chain. In particular, in Sec. V.1 we demonstrate that varying the gate voltage leads to a controlled SNS to SIS transition as the Fermi level enters the gaps of the Fibonacci chain. Then, we show in Sec. V.2 how the winding of the subgap states in the different gaps lead to distinct oscillations in the critical current ${I_{\mathrm{c}}}$, such that each winding number (i.e. gap label) can be determined, thus effectively measuring the topological invariant. This occurs when the Fermi level is located inside or close to the respective gap, thus accessible by tuning the applied gate voltage.
V.1 Gate-voltage tuning: fractal SNS to SIS transitions
So far, most of our work has centered on systematic model calculations to establish a fundamental theoretical understanding of the Josephson effect in quasiperiodic junctions. While some of the model parameters investigated may not be tunable in situ in an experiment once the device has been fabricated, our calculations still provide important fundamental understanding of the influence of the quasiperiodic modulation. Here, we study the influence of the chemical potential $\mu$, as envisioned by the application of a gate voltage to the non-superconducting part of the junction. Such a gate voltage is highly accessible in experiment, and our following results predict that it should lead to tunable and directly observable fractal oscillations between SNS and SIS behavior.
@@ V.1 Gate-voltage tuning: fractal SNS to SIS transitions
Next, we focus on the quasiperiodic scenario $\rho\neq 1$ and note the same features with peaks and overall larger current close to half-filling (but lower bandwidth $|\mu|<1.8{t_{\mathrm{B}}}$). Importantly, however, the quasiperiodic gap structure is directly probed by a sudden drop in current as the Fermi level enters the quasiperiodic gaps ${\Delta_{\mathrm{Q}}}$. We find that the size of these gaps relative to the proximitized superconducting gap $|\Delta|$ determines the junction behavior. In particular, inside the larger gaps (${\Delta_{\mathrm{Q}}}>|\Delta|$) we find SIS behavior with the current becoming orders of magnitude smaller than outside the gaps where there is instead SNS behavior. Inside the smaller gaps (${\Delta_{\mathrm{Q}}}<|\Delta|$) the behavior strongly depends on the junction length. For instance, in the shorter junction in Fig. 12(a) the current is still the same order of magnitude as outside the gaps due to pronounced ABS and supergap contributions, and thus of SNS type, but the gaps still cause a reduction or even complete absence of some peaks that are otherwise present in the crystalline junction. In contrast, in the longer junction in Fig. 12(b), the current is more strongly suppressed in the small gaps, showing more SIS-type behavior. Thus, the gap ratio ${\Delta_{\mathrm{Q}}}/|\Delta|$ together with the Fibonacci chain length can qualitatively alter the junction properties, and the SNS to SIS transitions in $\mu$ are fractal since the Fibonacci spectrum itself is fractal [17]. We note that the gap ratio was recently shown to also be important for both emergent topological superconductivity [76] and intrinsic superconductivity [73].
@@
Finally, we briefly comment on the overall magnitude of ${I_{\mathrm{c}}}$ in quasiperiodic versus crystalline junctions when varying ${t_{\mathrm{int}}}\in[0.1,1.0]{t_{\mathrm{B}}}$, $\rho\in[0.1,1.4]$ and $\mu$ between the band edges. Similar to Sec. IV.2, we find that the crystalline junction in most cases has a higher critical current due to the idealized perfect ballistic model with an exact zero-energy state. Only in a few parts of the parameter space does the quasiperiodic junction provide a significantly higher ${I_{\mathrm{c}}}$ than the crystalline junction, usually when ${t_{\mathrm{int}}}$ is closer to the quasiperiodically modified expression for ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, or when ${t_{\mathrm{int}}}$ is small and $\rho$ large.
@@
V.2 Phason angle dependence and topological invariant from critical current
In the previous Sec. V.1 we investigated the behavior of $\mu$ on ${I_{\mathrm{c}}}$, demonstrating that the junction can change between SNS and SIS behavior whenever the Fermi level enter the largest topological gaps of the Fibonacci chain. In this section, we show how the critical current can measure the winding number in each of these gaps. We begin by studying the influence of the phason angle $\phi$, in order to better understand and quantify how phason modes influence the Josephson physics. Similar to phonons, phason modes propagate through quasiperiodic materials where they induce phason flips [17], and we demonstrate that they have quantifiable influence on the critical current. We then describe the principle idea behind connecting the winding number to the critical current, followed by our results and proof-of-principle. Beyond providing a theoretical understanding of the influence of phason modes, these results could be measured e.g. by using STM techniques [102, 105, 117] to fabricate an ensemble of junctions each with a different phason angle, or exciting phason modes, or realizing an effective Fibonacci hopping model in a metamaterial [161, 162, 163, 164, 165] with quantum dots where phason flips are induced by in situ tuning the the coupling between the quantum dots.